if x varies directly as y and x = 6 when y = 15, find the value of x when y = 25
Answer by mananth(16949)

if x varies directly as y and x = 6 when y = 15, find the value of x when y = 25
xalphay
x=ky
...
plug value of x & y to get value of k
6=k15
k=2/5
...
x=ky
y=25, k = 2/5
x=2/5 * 25
x= 10
..